# David Martínez-Crespo1,2, Cesare Tronci3 1Departamento de Física Teórica, Universidad de Zaragoza, Zaragoza, Spain 2Centro de Astropartículas y Física de Altas Energías, Universidad de Zaragoza, Zaragoza, Spain

### v1

@@ 1.1 Models of hybrid quantum-classical dynamics
In recent years, our team has proposed an alternative model of mixed quantum-classical dynamics that not only retains physical consistency, but also reproduces decoherence levels with high accuracy. Successfully tested in [7] for several problems in physics and chemistry, the model is based on the theory of Koopman wavefuctions [33, 37] and builds up on early work by George Sudarshan [15, 42]. Koopman wavefunctions are square-integrable phase-space functions $\chi(q,p,t)$ such that the density $f=|\chi|^{2}$ satisfies the classical Liouville equation. In [42], Sudarshan proposed to consider mixed quantum-classical dynamics on the tensor-product Hilbert space of Koopman and Schrödinger wavefunctions. For example, in the case of infinite-dimensional quantum systems, this leads to writing $\widehat{P}(q,p,x,x^{\prime})=\Upsilon(q,p,x)\Upsilon(q,p,x^{\prime})^{*}$. Here, $\Upsilon(q,p,x)$ is a hybrid quantum-classical wavefunction, so that $f(q,p)=\int|\Upsilon(q,p,x)|^{2}\,{\rm d}x$ and we recall $\hat{\rho}=\widehat{P}/f$. While the construction of a dynamical model on such a tensor-product Hilbert space is a very intuitive idea, its implementation is far from obvious and only recently was our team able to present the resulting equations of motion. Their formulation went through several stages [10, 23, 24] and has been reviewed in [44]. Importantly, the quantum-classical model based on Koopman wavefunctions is Hamiltonian and, similarly to the Ehrenfest equations (1), enjoys a variational formulation in terms of Hamilton’s action principle. In particular, the model appears as an $\hbar$-correction of Ehrenfrest dynamics, where the correcting terms make the corresponding equations of motion rather formidable. The latter are as follows:
@@
^$\,\widehat{\!\mathscr{H}}=\widehat{H}+i\hbar\Big{(}\{\hat{\rho},{\widehat{H}}\}+\{\widehat{H},\hat{\rho}\}-\frac{1}{2}[\{\ln f,\widehat{H}\},\hat{\rho}]\Big{)}.$ (4)
Despite their intimidating appearance, it was pointed out in [7, 44] that these equations do not involve gradients of order higher than two. This is in contrast, for example, with fully quantum approaches based on quantum hydrodynamics that contain third-order derivatives [34]. Nevertheless, expanding the different terms in the phase-space vector field $\boldsymbol{X}$ makes this system hardly approachable by conventional methods. Luckily, however, the equations (2)-(4) enjoy a variational formulation which has so far served as the main operational platform for devising various extensions and reduced models, from quantum-classical spin dynamics [20] to fluid closures [22] and numerical particle schemes [7].
@@ 1.2 Mixed quantum-classical Heisenberg dynamics
So far, most works on mixed quantum-classical models have only dealt with the dynamics of states, i.e. the quantum and the classical states in interaction. A few authors, however, have posed the more fundamental question of the interaction of quantum and classical observables, as given by operator-valued functions on phase-space [6, 26, 38, 40, 14]. In particular, special emphasis has been put on how the quantum backreaction affects the classical motion. Also, questions emerge about the algebraic nature of these observables: given that classical and quantum observables comprise a Lie-algebra structure under the Poisson bracket and the commutator, respectively, is there an analogue Lie-algebra structure for quantum-classical observables? Similar questions were recently investigated in [24] by resorting to the geometry of hybrid quantum-classical wavefunctions. In this paper, we show that such a Lie-algebra structure may generally be absent.
@@
Following this introductory discussion, in Section 1.3 we review the Eulerian formulation of the Koopman hybrid model (2)-(4) based on Euler-Poincaré reduction by symmetry. Particular emphasis will be made on the appearance of the Wilczek-Zee connection in the Hamiltonian functional. As we will see, this connection plays a prominent role in quantum-classical Heisenberg dynamics. In Section 2 we deal with the Heisenberg picture of the Ehrenfest model. In spite of its simplicity, this model retains all the essential geometric properties that will also be present in the more advanced context of Koopman quantum-classical dynamics. We will discuss the purely classical and quantum specializations of the description along as well as comment on the conservation laws in the model. In Section 3 we use similar techniques to deal with the Koopman model. The classical part of the dynamics, i.e. the generator of diffeomorphic phase-space paths, deserves special attention and thus we devote Section 3.1 to its study. In particular we develop a new algebraic tool that, inside a pairing, acts over operators as a Lie derivative with respect to an operator-valued vector field. This remarkable tool allows us to unveil the geometric nature underlying the quantum back-reaction in the Koopman hybrid model. We use this approach in 3.2 to present the equations of motion of an arbitrary hybrid operator undergoing hybrid Koopman dynamics. Conservation laws are also discussed again in this more involved setting. In Section 4 we exemplify the Heisenberg equations of motion of the Koopman model for the simple case of a pure-dephasing Hamiltonian. Finally we conclude in Section 5 with a general discussion and we provide an outline of possible future directions.
@@ 1.3 Quantum-classical variational setting and gauge connections
The gradients appearing in the last term of (8) may be expressed in terms of a non-Abelian gauge connection first appeared in the work by Wilczek and Zee [46] and playing a predominant role in the main part of this paper. To see how this connection emerges, we can use the second in (6) to compute $\nabla\hat{\rho}=[\hat{\rho},\widehat{\boldsymbol{\lambda}}]$, where
@@ 2 Heisenberg picture of hybrid Ehrenfest dynamics
$\int\!f\big{(}\boldsymbol{\cal A}\cdot\boldsymbol{X}+\langle\hat{\rho},i\hbar\hat{\xi}-\boldsymbol{\cal X}\widehat{H}\rangle\big{)}\,{\rm d}^{2}z=\int\!f_{0}\big{(}\boldsymbol{\Theta}\cdot\boldsymbol{\cal X}+\langle\hat{\rho}_{0},i\hbar\widehat{\varpi}-\widehat{\cal H}\rangle\big{)}\,{\rm d}^{2}z_{0}$
@@
We observe that the Heisenberg Hamiltonian does not evolve on a group orbit because the map $\widehat{H}\mapsto{U}^{\dagger}(\eta^{*}\widehat{H}){U}$ does not correspond to a group action. In order to restore the information on group orbits, here we will express the unitary operator $U$ in (6) in the classical convective frame by writing $U={\cal U}\circ\boldsymbol{\eta}$, without loss of generality. Then, we obtain $\dot{U}=\partial_{t}{\cal U}\circ\boldsymbol{\eta}+{\boldsymbol{\cal X}}\cdot\nabla({\cal U}\circ\boldsymbol{\eta})$ and, by a slight abuse of notation, the forth equation in (11) becomes
@@
We observe that the dynamics has now moved from the state variables $f$ and $\hat{\rho}$ to the Hamiltonian operator $\widehat{\cal H}$ as well as the auxiliary quantities $\Theta=\boldsymbol{\Theta}\cdot{\rm d}{\mathbf{z}}$ and $\widehat{{\gamma}}=\widehat{\boldsymbol{\gamma}}\cdot{\rm d}{\mathbf{z}}$. As we will see, the Lie-transport of the canonical one-form means that the symplectic properties of the classical phase-space are now evolved in time. By taking the relevant derivatives in (11)-(12), we find the auxiliary equations
@@
respectively. Here, the notation $\nabla^{\widehat{{\gamma}}\,}\cdot=\nabla\cdot\,+\,[\widehat{\boldsymbol{\gamma}},\cdot\,]$ is used whenever we work in local coordinates. More importantly, the variations in (14) are accompanied by the relations
@@
The arbitrary skew-Hermitian operator $\hat{\zeta}_{0}$ commuting with the initial state $\hat{\rho}_{0}$ may appear as a new unexpected object in the present construction of Heisenberg dynamics. However, this arbitrariness is unavoidable and far from accidental. This may be easily seen by considering the second evolution law in (6), that is ${\hat{\rho}=(U\hat{\rho}_{0}U^{\dagger})\circ\boldsymbol{\eta}^{-1}}$. It is clear that the unitary operator $U$ is only defined up to right-multiplication by any unitary operator $U_{0}$ preserving the initial state $\hat{\rho}_{0}$, that is $U_{0}\hat{\rho}_{0}U_{0}^{\dagger}=\hat{\rho}_{0}$. In other words, the propagator $U$ is only defined up to elements of the isotropy group $G_{0}=\{U_{0}\in{\cal F}(T^{*}Q,\mathscr{U}(\mathscr{H}_{\text{\tiny Q}}))\,|\,U_{0}\hat{\rho}_{0}U_{0}^{\dagger}=\hat{\rho}_{0}\}$. This corresponds to a non-Abelian gauge freedom also appearing in the standard quantum treatment of Heisenberg dynamics, as already discussed in [12]. Accordingly, the infinitesimal generator $\hat{\zeta}$ of the quantum motion is defined up to elements of the isotropy subalgebra $\mathfrak{g}_{0}=\{\hat{\zeta}_{0}\in{\cal F}(T^{*}Q,\mathfrak{u}(\mathscr{H}_{\text{\tiny Q}}))\,|\,[\hat{\rho}_{0},\hat{\zeta}_{0}]=0\}$. Since this skew-hermitian operator $\hat{\zeta}_{0}$ is an arbitrary gauge choice, one may safely set $\hat{\zeta}_{0}=0$ in the final equations. This is the gauge fixing performed throughout this paper.
@@
where ${\mathcal{A}}_{B,0}=\langle\psi_{0},-i\hbar{\rm d}\psi_{0}\rangle$ and we recall ${\cal U}=\eta_{*}U$. Thus, if we introduce the Berry curvature ${\mathcal{B}}={\rm d}{\mathcal{A}}_{B}$ and recall $\Omega=\eta^{*}\omega$, taking the exterior differential of (19) yields
@@
which recovers the relation ${\omega+{\mathcal{B}}=\eta_{*}(\omega+{\mathcal{B}}_{0})}$ found in the Schrödinger picture [21, 23].
@@
In the purely classical case, we have ^$\widehat{\cal H}={\cal H}\boldsymbol{1}$ and $\widehat{\boldsymbol{\gamma}}={\boldsymbol{\gamma}}\boldsymbol{1}$, so that $\partial_{t}\Omega=0$ and ${\cal H}=H$. We recognize the usual evolution ${\partial_{t}{{\cal O}}}=\omega^{-1}({\rm d}{{\cal O}},{{\rm d}{{H}}})=\{{{\cal O}},{H}\}$ for the classical observable $\widehat{{\cal O}}={{\cal O}}\boldsymbol{1}$. Once again, the purely quantum case is recovered by setting ${\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}}=0$.
Remark 4 (Conservation laws)
@@
Here we can distinguish three cases. In the first case $\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})=0$, one says that $\mathcal{X}_{\widehat{\mathcal{O}}}$ and $\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle$ are symplectically orthogonal and the equation of motion reduces to the Heisenberg equation
$\frac{\partial\widehat{{\cal O}}}{\partial t}=i\hbar^{-1}[{\widehat{\mathcal{H}}},\widehat{{\cal O}}].$
The symplectic orthogonality condition is written in the Eulerian frame as $0=\omega({X}_{\widehat{{O}}},{{\langle{X}_{\widehat{{H}}}\rangle}})=\pounds_{\langle X_{\widehat{{H}}}\rangle}{\widehat{{O}}}$. Such an observable is unaffected by the action of the diffeomorphism in the evolution. Therefore, if the observable is purely classical, it appears as a constant of motion. In the second case, we have [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}]=0$, so that
$\frac{\partial\widehat{{\cal O}}}{\partial t}=\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}}).$
In this case the apparent quantum evolution of the observable is determined by the presence of the connection $\widehat{\gamma}$ in the construction of the vector fields. The appearance of $\widehat{\gamma}$, however, is simply related with the gauge nature of the quantum frame induced by $\mathcal{U}$ and should not be understood as a proper quantum evolution. When this type of observable is purely quantum, one has ${\rm d}^{\widehat{\gamma}}\widehat{\mathcal{O}}=0$, so that $\widehat{\mathcal{O}}$ appears as a constant of motion. In the third case, both $\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})=0$ and [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}]=0$, so that the whole operator $\widehat{{\cal O}}$ is conserved. As we saw before, a more likely possibility is that both ${\Omega(\langle\mathcal{X}_{\widehat{\mathcal{O}}}\rangle,{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})=0}$ and [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}]=0$ hold, in which case the local expectation $\langle{\widehat{\mathcal{O}}}\rangle$ is conserved.

### v2

@@ 1.1 Models of hybrid quantum-classical dynamics
In recent years, our team has proposed an alternative model of mixed quantum-classical dynamics that not only retains physical consistency, but also reproduces decoherence levels with high accuracy. Successfully tested in [7] for several problems in physics and chemistry, the model is based on the theory of Koopman wavefuctions [37, 41] and builds up on early work by George Sudarshan [16, 46]. Koopman wavefunctions are square-integrable phase-space functions $\chi(q,p,t)$ such that the density $f=|\chi|^{2}$ satisfies the classical Liouville equation. In [46], Sudarshan proposed to consider mixed quantum-classical dynamics on the tensor-product Hilbert space of Koopman and Schrödinger wavefunctions. For example, in the case of infinite-dimensional quantum systems, this leads to writing $\widehat{P}(q,p,x,x^{\prime})=\Upsilon(q,p,x)\Upsilon(q,p,x^{\prime})^{*}$. Here, $\Upsilon(q,p,x)$ is a hybrid quantum-classical wavefunction, so that $f(q,p)=\int|\Upsilon(q,p,x)|^{2}\,{\rm d}x$ and we recall $\hat{\rho}=\widehat{P}/f$. While the construction of a dynamical model on such a tensor-product Hilbert space is a very intuitive idea, its realization is far from obvious and only recently was our team able to present the resulting equations of motion. Their formulation went through several stages [11, 26, 27] and has been reviewed in [49]. Importantly, the quantum-classical model based on Koopman wavefunctions is Hamiltonian and, similarly to the Ehrenfest equations (1), enjoys a variational formulation in terms of Hamilton’s action principle. In particular, the model appears as an $\hbar$-correction of Ehrenfrest dynamics, where the correcting terms make the corresponding equations of motion rather formidable. The latter are as follows:
@@
$\,\widehat{\!\mathscr{H}}$ $=\widehat{H}+i\hbar\Big{(}\{\hat{\rho},{\widehat{H}}\}+\{\widehat{H},\hat{\rho}\}-\frac{1}{2}[\{{\ln}f,\widehat{H}\},\hat{\rho}]\Big{)},$ (4)
Despite their intimidating appearance, it was pointed out in [7, 49] that these equations do not involve gradients of order higher than two. This is in contrast, for example, with fully quantum approaches based on quantum hydrodynamics that contain third-order derivatives [38]. Nevertheless, expanding the different terms in the phase-space vector field $\boldsymbol{X}$ makes this system hardly approachable by conventional methods. Luckily, however, the equations (2)-(4) enjoy a variational formulation which has so far served as the main operational platform for devising various extensions and reduced models, from quantum-classical spin dynamics [23] to fluid closures [25] and numerical particle schemes [7]. Recently, we also identified an entropy functional for this system [48].
@@ 1.2 Mixed quantum-classical Heisenberg dynamics
So far, most works on mixed quantum-classical models have only dealt with the dynamics of states, i.e. the quantum and the classical states in interaction. A few authors, however, have posed the more fundamental question of the interaction of quantum and classical observables, as given by operator-valued functions on phase-space [6, 29, 42, 44, 15]. In particular, special emphasis has been put on how the quantum backreaction affects the classical motion. Also, questions emerge about the algebraic nature of these observables: given that classical and quantum observables comprise a Lie-algebra structure under the Poisson bracket and the commutator, respectively, is there an analogue Lie-algebra structure for quantum-classical observables? Similar questions were recently investigated in [27] by resorting to the geometry of hybrid quantum-classical wavefunctions. While this paper shows how the evolution of hybrid observables is generated by a Lie-algebra action underlying a group structure, the observables themselves fail to inherit such a Lie-algebra structure.
@@
Following this introductory discussion, in Section 1.3 we review the Eulerian formulation of the Koopman hybrid model (2)-(4) based on Euler-Poincaré reduction by symmetry. Particular emphasis will be placed on the appearance of the Wilczek-Zee connection in the Hamiltonian functional. As we will see, this connection form plays a prominent role in quantum-classical Heisenberg dynamics. In Section 2 we deal with the Heisenberg picture of the Ehrenfest model. In spite of its simplicity, this model retains all the essential geometric properties that will also be present in the more advanced context of Koopman quantum-classical dynamics. We will discuss the purely classical and quantum specializations of the description along as well as comment on the conservation laws in the model. In Section 3 we use similar techniques to deal with the Koopman model. The classical part of the dynamics, i.e. the generator of diffeomorphic phase-space paths, deserves special attention and thus we devote Section 3.1 to its study. In particular we develop a new algebraic tool that, inside a pairing, acts over operators as a Lie derivative with respect to an operator-valued vector field. This valuable tool allows us to unveil the geometric nature underlying the quantum back-reaction in the Koopman hybrid model. We use this approach in 3.2 to present the equations of motion of an arbitrary hybrid operator undergoing hybrid Koopman dynamics. Conservation laws are also discussed again in this more involved setting. In Section 4 we exemplify the Heisenberg equations of motion of the Koopman model for the simple case of a pure-dephasing Hamiltonian. Finally we conclude in Section 5 with a general discussion and we provide an outline of possible future directions.
@@ 1.3 Quantum-classical variational setting and gauge connections
The gradients appearing in the last term of (8) may be expressed in terms of a non-Abelian gauge connection first appeared in the work by Wilczek and Zee [51] and playing a predominant role in the main part of this paper. To see how this connection emerges, we can use the second equation in (6) to compute $\nabla\hat{\rho}=[\hat{\rho},\widehat{\boldsymbol{\lambda}}]$, where
@@ 2 Heisenberg picture of hybrid Ehrenfest dynamics
$\int\!f\big{(}\boldsymbol{\cal A}\cdot\boldsymbol{X}+\langle\hat{\rho},i\hbar\hat{\xi}-\widehat{H}\rangle\big{)}\,{\rm d}^{2}z=\int\!f_{0}\big{(}\boldsymbol{\Theta}\cdot\boldsymbol{\cal X}+\langle\hat{\rho}_{0},i\hbar\widehat{\varpi}-\widehat{\cal H}\rangle\big{)}\,{\rm d}^{2}z_{0}$
@@
We observe that the Heisenberg Hamiltonian does not evolve on a group orbit because the map $\widehat{H}\mapsto{U}^{\dagger}(\eta^{*}\widehat{H}){U}$ does not correspond to a group action. In order to restore the information on group orbits, here we will express the unitary operator $U$ in (6) in the classical convective frame by writing $U={\cal U}\circ\boldsymbol{\eta}$, without loss of generality. Here, the initial conditions are such that $U|_{t=0}=\boldsymbol{\eta}|_{t=0}=\boldsymbol{1}$. Then, also ${\cal U}|_{t=0}=\boldsymbol{1}$ and we obtain $\dot{U}=\partial_{t}{\cal U}\circ\boldsymbol{\eta}+{\boldsymbol{\cal X}}\cdot\nabla({\cal U}\circ\boldsymbol{\eta})$. By a slight abuse of notation, the forth equation in (11) becomes
@@
We observe that the dynamics has now moved from the state variables $f$ and $\hat{\rho}$ to the Hamiltonian operator $\widehat{\cal H}$ as well as the auxiliary quantities $\Theta=\boldsymbol{\Theta}\cdot{\rm d}{\mathbf{z}}$ and $\widehat{{\gamma}}=\widehat{\boldsymbol{\gamma}}\cdot{\rm d}{\mathbf{z}}$. As we will see, the presence of the Lie-transported canonical one-form $\Theta=\eta^{*}{\cal A}$ means that the symplectic properties of the classical phase-space are now evolved in time. By taking the relevant derivatives in (11)-(12), we find the auxiliary equations
@@
respectively. Here, the notation $\nabla^{\widehat{{\gamma}}\,}\cdot=\nabla\cdot\,+\,[\widehat{\boldsymbol{\gamma}},\cdot\,]$ is used whenever we work in local coordinates.
Remark 2 (Evolution of the connection form)
[…]
More importantly, the variations in (14) are accompanied by the relations
@@
The arbitrary skew-Hermitian operator $\hat{\zeta}_{0}$ commuting with the initial state $\hat{\rho}_{0}$ may appear as a new unexpected object in the present construction of Heisenberg dynamics. However, this arbitrariness is unavoidable and far from accidental. This may be easily seen by considering the second evolution law in (6), that is ${\hat{\rho}=(U\hat{\rho}_{0}U^{\dagger})\circ\boldsymbol{\eta}^{-1}}$. It is clear that the unitary operator $U$ is only defined up to right-multiplication by any unitary operator $U_{0}$ preserving the initial state $\hat{\rho}_{0}$, that is $U_{0}\hat{\rho}_{0}U_{0}^{\dagger}=\hat{\rho}_{0}$. In other words, the propagator $U$ is only defined up to elements of the isotropy group $G_{0}=\{U_{0}\in{\cal F}(T^{*}Q,\mathscr{U}(\mathscr{H}_{\text{\tiny\sf Q}}))\,|\,U_{0}\hat{\rho}_{0}U_{0}^{\dagger}=\hat{\rho}_{0}\}$. This corresponds to a non-Abelian gauge freedom also appearing in the standard quantum treatment of Heisenberg dynamics, as already discussed in [13]. Accordingly, the infinitesimal generator $\hat{\zeta}$ of the quantum motion is defined up to elements of the isotropy subalgebra $\mathfrak{g}_{0}=\{\hat{\zeta}_{0}\in{\cal F}(T^{*}Q,\mathfrak{u}(\mathscr{H}_{\text{\tiny\sf Q}}))\,|\,[\hat{\rho}_{0},\hat{\zeta}_{0}]=0\}$. Since this skew-hermitian operator $\hat{\zeta}_{0}$ is an arbitrary gauge choice, one may safely set $\hat{\zeta}_{0}=0$ in the final equations. This is the gauge fixing performed throughout this paper.
@@
where ${\mathcal{A}}_{B,0}=\langle\psi_{0},-i\hbar{\rm d}\psi_{0}\rangle$ and we recall ${\cal U}=\eta_{*}U$. Thus, if we introduce the Berry curvature ${\mathcal{B}}={\rm d}{\mathcal{A}}_{B}$ and denote $\Omega=\eta^{*}\omega$, taking the exterior differential of (19) yields
@@
which recovers the relation ${\omega{+}{\mathcal{B}}=\eta_{*}(\omega{+}{\mathcal{B}}_{0})}$ already found in the Schrödinger picture [24, 26].
@@
In the purely classical case, we have ^$\widehat{\cal H}={\cal H}\boldsymbol{1}$ and $\widehat{\boldsymbol{\gamma}}={\boldsymbol{\gamma}}\boldsymbol{1}$, so that $\partial_{t}\Omega=0$ and ${\cal H}=H$. We recognize the usual evolution ${\partial_{t}{{\cal O}}}=\omega^{-1}({\rm d}{{\cal O}},{{\rm d}{{H}}})=\{{{\cal O}},{H}\}$ for the classical observable $\widehat{{\cal O}}={{\cal O}}\boldsymbol{1}$. Once again, the purely quantum case is recovered by setting ${\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}}=0$. We emphasize that the local expectation values $\langle\widehat{O}\rangle$ must not be confused with the overall expectation values ${\int}f_{0}\langle\widehat{O}\rangle{\rm d}^{2}z_{0}$, which are the measurable quantities. As we will see later, the overall expectation may be conserved even if the local expectation is not.
Remark 5 (Special cases and conservation laws)
@@
Here we can distinguish three special cases. In the first case, $\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})=0$, one says that $\mathcal{X}_{\widehat{\mathcal{O}}}$ and $\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle$ are symplectically orthogonal and the equation of motion reduces to the Heisenberg equation ${\partial_{t}\widehat{{\cal O}}}=i\hbar^{-1}[{\widehat{\mathcal{H}}},\widehat{{\cal O}}]$. The symplectic orthogonality condition is written in the Eulerian frame as $0=\omega({X}_{\widehat{{O}}},{{\langle{X}_{\widehat{{H}}}\rangle}})=\pounds_{\langle X_{\widehat{{H}}}\rangle}{\widehat{{O}}}$. Such an observable is unaffected by the action of the diffeomorphism in the evolution. This is the case for a purely quantum observable $\widehat{\cal O}=\eta^{*}({\cal U}^{\dagger}\widehat{O}{\cal U})$. In the second case, we have [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}]=0$, so that $^{\partial_{t}\widehat{{\cal O}}}=\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})$. In this case the apparent quantum evolution of the observable is determined by the presence of the connection $\widehat{\gamma}$ in the construction of the vector fields. The appearance of $\widehat{\gamma}$, however, is simply related with the gauge nature of the quantum frame induced by $\mathcal{U}$ and should not be understood as a proper quantum evolution. This is the case for a purely classical observable. In the third case, both $\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})=0$ and [^$[{\widehat{\mathcal{H}}},\widehat{{\cal O}}]=0$, so that the operator $\widehat{{\cal O}}$ is conserved.
One can also consider the conservation of local expectations, which may result from the combination of different properties. For example, the local expectation $\langle\widehat{\cal N}\rangle$ of the quantity $\widehat{\cal N}=({\mathcal{P}}^{2}+{\mathcal{Q}}^{2}+\widehat{{\Sigma}}_{z})/2$ is conserved by the Jaynes-Cummings Hamiltonian ^$\widehat{\mathcal{H}}=\widehat{\cal N}+g({\mathcal{Q}}\widehat{{\Sigma}}_{x}-{\mathcal{P}}\widehat{{\Sigma}}_{y})$, where $({\cal Q},{\cal P})=(\eta^{*}q,\eta^{*}p)$ and, in the usual Pauli matrix notation, $\widehat{\boldsymbol{\Sigma}}=\eta^{*}({\cal U}^{\dagger}\widehat{\boldsymbol{\sigma}}{\cal U})$. This conservation is verified directly upon using $0=\eta^{*}({\cal U}^{\dagger}{\rm d}\widehat{\boldsymbol{\sigma}}{\cal U})={\rm d}^{\widehat{\gamma}}\widehat{\boldsymbol{\Sigma}}$ so that, for example, $\mathcal{X}_{{\mathcal{Q}}\widehat{{\Sigma}}_{x}}=\widehat{{\Sigma}}_{x}\mathcal{X}_{{\mathcal{Q}}}$. Similarly, one shows the conservation of the expectation $\langle\widehat{\cal M}\rangle$ of the total momentum ^$\widehat{\cal M}={\mathcal{P}}{+}\,\widehat{\!\mathscr{K}}$ for translation invariant Hamiltonians of the type ^$\widehat{\mathcal{H}}=(M^{-1}{\mathcal{P}}^{2}+m^{-1}\widehat{\!\mathscr{K}\,}\!^{2})/2+V(\mathcal{Q}-\widehat{\!\mathscr{X}})$, where $(\widehat{\!\mathscr{X}},\widehat{\!\mathscr{K}})=\eta^{*}({\cal U}^{\dagger}(\widehat{x},\widehat{p}){\cal U})$ are quantum canonical observables such that $[\widehat{\!\mathscr{X}},\widehat{\!\mathscr{K}}]=i\hbar$. Here, the conservation of $\langle\widehat{\cal M}\rangle$ follows from ${\rm d}^{\widehat{{\gamma}}\,}\widehat{\!\mathscr{X}}=0$, so that ${\rm d}^{\widehat{{\gamma}}}V=\partial_{\mathcal{Q}}V{\rm d}\mathcal{Q}$ and $\Omega(\mathcal{X}_{\widehat{\mathcal{O}}},{{\langle\mathcal{X}_{\widehat{\mathcal{H}}}\rangle}})=\Pi({\rm d}{{\cal P}},{\langle{\rm d}^{\widehat{{\gamma}}}\widehat{\mathcal{H}}\rangle})=\hbar^{-1}\langle i[\,\widehat{\!\mathscr{K}},{\widehat{\mathcal{H}}}]\rangle$.